OpenStudy (anonymous):

can some one show me step by step of how to simplify the square root of 18

OpenStudy (anonymous):

\[\sqrt{18} = \sqrt{9*2} = \sqrt{3*3*2} \] as 9 is square of 3 then square root of 9 is 3 so \[\sqrt{18} = 3\sqrt{2}\]

OpenStudy (anonymous):

i still dont get how u got yo anwser

OpenStudy (jdoe0001):

\(\large \bf { \sqrt{18}\qquad \begin{cases} 18\to 3\cdot 3\cdot 2\\ 18\to 3^2\cdot 2 \end{cases}\implies \sqrt{18}\implies \sqrt{3^2\cdot 2} \\ \quad \\ \sqrt[{\color{brown}{ 2}}]{3^{\color{brown}{ 2}}}\cdot \sqrt[{\color{brown}{ 2}}]{2}\implies 3\cdot \sqrt[{\color{brown}{ 2}}]{2}\implies 3\sqrt{2} }\) see it now?

OpenStudy (anonymous):

which part?

OpenStudy (anonymous):

the anwser

OpenStudy (anonymous):

@jdoe0001 explanation worked?

OpenStudy (anonymous):

i got it